OpenStudy (anonymous):

The cost to produce a product is modeled by the function f(x) = 5x^2 − 70x + 258 where x is the number of products produced. Complete the square to determine the minimum cost of producing this product. A. (x − 7)^2 + 13; The minimum cost to produce the product is $13. B. 5(x − 7)^2 + 13; The minimum cost to produce the product is $7. C. 5(x − 7)^2 + 258; The minimum cost to produce the product is $7. D. 5(x − 7)^2 + 258; The minimum cost to produce the product is $258.

OpenStudy (mertsj):

\[5x^2-70x+258=5(x^2-14x+.......)+258\]

OpenStudy (mertsj):

\[5(x^2-14x+49)+258-245= 5(x-7)^2+13\]